MEDICATION CALCULATION #57

Order: aminocaproic acid 4 g in 250 mL D5W to infuse in 1 hour. Set: secondary set labeled 15 gtt/min. Available: (aminocaproic acid) Injection USP Intravenous 250 mg/mL

a. How many mL of aminocaproic acid should be mixed in 250 mL of D5W?

b. How many gtt/min should the client receive?

Please show your step by step calculations.
Answer #1

A)

Order follows 4 gm of a.a. in 250 ml of D5W.

1 ml of injection of a.a. contains 250 mg of a.a

So, a.a. injection to be mixed = 4gm/ 250 mg =4000mg/ 250 mg = 16 ml

Hence, 16 ml of aminocaproic acid should be mixed with 250mL of D5W

B)

250 mL of D5W to be mixed with16 mL of aminocaproic acid

So, (250+16) mL = 266 mL of solution is to be infused over 1 hour or 60 minutes.

Drop factor= 15 gtt/min

Drop calculation formula=

(Volume/Time ) ×Drop factor

= (266/60)×15 = 66.5 gtt/min

Hence, the client should receive 66.6 gtt/min